OpenStudy (anonymous):

The point (0, -3.5) lies on the line 11x – 6y = 21. Determine if this point is a solution of the system 11x – 6y = 21, 9x + 8y = 43

OpenStudy (chaise):

Simply substitute your respective x and y values into the equation in order to determine the solution. 6*0-6*-3.5=21 -6*-3.5=21 In your case, x is 0. and y is negative, in order to yield a positive answer, the coefficient infront of y must also be negative (two negatives make a positive) you can automatically infer, without any working out that the answer is the first option.

OpenStudy (anonymous):

Thanks! I can see the two negatives making the positive results.